# SIFU: Side-view Conditioned Implicit Function for Real-world Usable Clothed Human Reconstruction

### 23 Dec 2023 v2

@@ 2 Related Work
Explicit-shape-based Reconstruction. Recovering the human mesh from a single RGB image is a complex challenge that has received extensive attention. Many approaches [89, 16, 47, 45, 39, 41, 42, 90, 46, 40, 43, 17] adopt parametric body models [54, 84, 62, 38] to estimate the shape and pose of a 3D human body with minimal clothing. To incorporate clothing into the 3D models, methods often apply 3D clothing offsets [5, 2, 3, 4, 93, 44, 81] or use adjustable garment templates [8, 34, 17] over the base body shape. Additionally, non-parametric forms like depth maps [73, 19], normal maps [83], and point clouds [88] are explored for creating representations of clothed humans.
@@ 3.1 Preliminary
Diffusion Models. Diffusion processes, notably represented by Diffusion Probabilistic Models (DPM) [74, 28, 13, 12, 59], are pivotal in image generation. These models aim to approximate a data distribution $q$ through a progressive denoising process. Starting with a Gaussian i.i.d noisy image $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the model denoises it until a clean image $\bm{x}_{0}$ from the target distribution $q$ is obtained. DPMs can also learn a conditional distribution with additional guiding signals like text conditioning.
@@ 3.2 Side-view Conditioned Implicit Function
Pipeline. For a given input image and its reconstructed mesh ${M}$, we first utilize image-to-text models (such as GPT-4v [60]) to convert the image into a textual description ${P}$, and then back-project the mesh color onto a UV texture map ${{T}}$, following the approach in [76]. To visualize unseen mesh areas, differentiable rendering $\mathcal{I}$ is employed on mesh $M$, generating images of these invisible views:
@@
where $\bm{J}=\{J^{1},...,J^{n}\}$ corresponds to the refined views of $\bm{I}$. After obtaining $\bm{J}$, a pixel-wise Mean Squared Error (MSE) loss is computed between each $J^{i}$ and $I^{i}$ to optimize the texture map $T$. Additional losses include a perceptual loss $\mathcal{L}_{vgg}$ [37] and a Chamfer Distance loss $\mathcal{L}_{CD}$ [31], aimed at ensuring style similarity between $\mathcal{J}$ and the input image. We also compute an MSE loss $\mathcal{L}^{f}_{MSE}$ from the input view against the input image. These combined losses jointly optimize $T$, enhancing overall texture quality:
@@ 4.1 Evaluation
For texture reconstruction, we base our comparisons on results from [11, 91, 1] due to the limited public availability of some models. Our tests on THuman2.0 [86], utilizing PIXIE [16] for SMPL-X estimations, ensure fair benchmarking. As shown in Fig. 5, SIFU surpasses PIFu [71] by 22.2% in PSNR, demonstrating its superior texture quality. For visual comparisons, refer to Fig. 5 and the SupMat.

### 08 Apr 2024 v3

@@ 2 Related Work
Explicit-shape-based Reconstruction. Recovering the human mesh from a single RGB image is a complex challenge that has received extensive attention. Many approaches [95, 16, 49, 46, 40, 42, 43, 96, 47, 41, 44, 17] adopt parametric body models [56, 88, 65, 39] to estimate the shape and pose of a 3D human body with minimal clothing [76, 48]. To incorporate clothing into the 3D models, methods often apply 3D clothing offsets [5, 2, 3, 4, 99, 45, 85] or use adjustable garment templates [8, 35, 17] over the base body shape. Additionally, non-parametric forms like depth maps [77, 19], normal maps [87], and point clouds [94] are explored for creating representations of clothed humans.
@@ 3.1 Preliminary
Diffusion Models. Diffusion processes, notably represented by Diffusion Probabilistic Models (DPM) [78, 28, 13, 12, 62], are pivotal in image generation and have shown capabilities in human/avatar generation [30, 89]. These models aim to approximate a data distribution $q$ through a progressive denoising process. Starting with a Gaussian i.i.d noisy image $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the model denoises it until a clean image $\bm{x}_{0}$ from the target distribution $q$ is obtained. DPMs can also learn a conditional distribution with additional guiding signals like text conditioning.
@@ 3.2 Side-view Conditioned Implicit Function
Pipeline. For a given input image and its reconstructed mesh ${M}$, we first utilize vision-to-text models (e.g.[63, 91, 58]) to convert the image into a textual description ${P}$, and then back-project the mesh color onto a UV texture map ${{T}}$, following the approach in [80]. To visualize unseen mesh areas, differentiable rendering $\mathcal{I}$ is employed on mesh $M$, generating images of these invisible views:
@@
where $\bm{J}=\{J^{1},...,J^{n}\}$ corresponds to the refined views of $\bm{I}$. After obtaining $\bm{J}$, a pixel-wise Mean Squared Error (MSE) loss is computed between each $J^{i}$ and $I^{i}$ to optimize the texture map $T$. Additional losses include a perceptual loss $\mathcal{L}_{vgg}$ [38] and a Chamfer Distance loss $\mathcal{L}_{CD}$ [32], aimed at ensuring style similarity between $\bm{I}$ and the input image. We also compute an MSE loss $\mathcal{L}^{f}_{MSE}$ from the input view against the input image. These combined losses jointly optimize $T$, enhancing overall texture quality:
@@ 4.1 Evaluation
For texture reconstruction, SIFU surpasses PIFu [74] by 22.2% in PSNR, demonstrating its superior texture quality. For visual comparisons, refer to Fig. 5 and the SupMat.
@@ 5 Conclusion
Acknowledgements. This work was supported by the National Natural Science Foundation of China (U2336212) and the Fundamental Research Funds for the Central Universities (No. 226-2022-00051).